N1. Find the wind direction and speed, given the (U, V) components a. (-5, 0) knots c. (-1,15) mi/h e. (8, 0) knots g. (-2,-10) mi/h b. (8,-2) m/s d. (6,6) m/s f. (5, 20) m/s h. (3,-3) m/s
Answer #1

a)

u = - 5 and v = 0

speed = sqrt(u2 + v2)

speed = sqrt((-5)2 + (0)2) = 5 knots

direction : tan-1(v/u) = tan-1(0/(-5)) = 180 deg from positive x-axis or east direction .

b)

u = 8 and v = - 2

speed = sqrt(u2 + v2)

speed = sqrt((8)2 + (- 2)2) = 8.25 m/s

direction : tan-1(v/u) = tan-1(-2/8) = - 14.04 deg

14.04 deg south of east

c)

u = - 1 and v = 15

speed = sqrt(u2 + v2)

speed = sqrt((- 1)2 + (15)2) = 15.03 mi/h

direction : tan-1(v/u) = tan-1(15/(-1)) = 93.8 deg counterclockwise from positive x-axis or east direction .

d)

u = 6 and v = 6

speed = sqrt(u2 + v2)

speed = sqrt((6)2 + (6)2) = 8.5 knots

direction : tan-1(v/u) = tan-1(6/6) = 45 deg counterclockwise from positive x-axis or east direction .